APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Thursday December 29, 2005 (MJD 53734) ---=== OBSERVING TEAM ===--- Swing: Dan Long Night: John Barentine Elena Malanushenko Viktor Malanushenko ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- We did 5 plates declared as "DONE" and collected signal on other 2 plates. Most of time we hade clear sky, but one time we stopped observations because clouds. We had one problem with rotator, which did not go to zero position after InstChange command. We moved rotator to zero before the command. (see log and problems). ---=== OBSERVING LOG ===--- Afternoon: No problems during the afternoon checkout. Spectrograph focus checked but not changed. Still set for 3 to 4 C. Details in problem section. 23:10Z Doors Open, Louvers Open, Fans On 00:15Z Enclosure Open Night: Start with spectroscopy as the weather is good but not perfect, to see the dynamic. Card 2, Plate 1542. GSOGTF. Twilight. 1:05 UT start observations. We did 4 exposures (3600 sec). We collected the signal r1=20.1 & r2=17.8, it was r1=12.0 & r2=12.1 already [MJD 53733]. Using idl>apoplotsky we got unusual strong spectral feature (10000 counts) around 6000A on 36009 data file. It has less intensity (8000) on 36010, and it looks as usual (5000) on 39011. Propably. we observed some atmospheric twilight line. We did pos-calibration as recommended us by SoS. "DONE". 2:49 UT Card 1, plate 1453. We have only 40 min to observe this plate, but the next plate starts in 40 minuts only. GSOGTF. We did 2 exposures (1800 sec) and collected s/n=5.2. "CONTINUE" Card 8, plate 2067. seguefaint. GSOGTF. Seguefaint. 3:40. We did 5 exposures (4500 sec) and collected some signal. As the time is expared we change plate to another. "CONTINUE" Problem: After command InstChange, the rotator stopped at 35 deg. After "axis init" and a track command the Rot stopped again with an error message about an overcurrent in the MCP menu. We did several attempts. Checked slips on the rack - they were ok, but we resert them anyway. Tried to move rotator to bigger angle, it worked. After several small steps to bigger angle we sent rotator to zero position, and it has been done. We did not understand what was a problem. We noticed in the MCP window when the rotator stopped that it was actually an overcurrent ("*AO*"), and also got two watcher errors; first, AxisNoSlew:rot (axis cannot slew: 2005-12-30 05:19:22Z) then, start move failed; offsetting and trying again (2005-12-30 05:27:00Z sdssmcp) (start_move failed; offsetting and trying again (2005-12-30 05:27:00Z sdssmcp): 2005-12-30 05:27:06Z ). The first error happened at approximately the time we sent the telescope to instChange. The second happened while we were at the telescope troubleshooting the problem. Also, some time during our attemptions we see large numbers at the error column of Rot on MCP GUI. Cart 7, plate 2051. 5:40 UT. Could not start observations because of clouds. The plate is expired before the sky is cleared to start the observation again. We have to replace the cartridge. "CONTINUE" InstChange got the same problem with Rot over-current at similar angle 33 deg. We used our previous recipe - move Rot to greater angle and after that move it to zero. It worked. We had error message when we had that problem: "Inconsistent instrument ID switches: 4 5 5 (2005-12-30 07:17:43Z sdssmcp) (Inconsistent instrument ID switches: 4 5 5 (2005-12-30 07:17:43Z sdssmcp): 2005-12-30 07:18:01Z )" 7:36 UT Cart 5, plate 2295. Started observations between clouds. Found Guider stars using fk5 star. Chunk. We did three exposures (2880 sec) and collected S/N=16.6. "DONE" The weather is better, no clouds, but the fwhm is around 2.0 and wind is stronger. To avoide the problem with rotator we made command tcc> rot 0 mount. before InstChange. 8:51 UT. Cart 4, plate 2349. Chunk. GSOGTF Previous s/n=6.5, observed on MJD 53731 and 53733. We did two more exposures (1980 sec) and got s/n=9.0. "DONE". 9:50 UT. Cart 9, plate 2346. Chunk. GSOGTF. We did three exposures (3300 sec) and got s/n=13.7, but all others are larger than 15. "DONE". 11:14 UT. Cart 3, plate 1975. Chunk. GSOGTF. We did threee exposures (4800 sec), got s/n=16.7. "DONE" endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 5962 21:20Z 21:36Z 100 O 127.76 131.54 38 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): yes QA Procedures Done (y/n): yes UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 35999, plate -9999 ------- 21:21 35999 0.0 bias ----- sequence 36000, plate 1542 ------- 21:25 36000 10.0 flat 21:27 36001 2.0 arc 21:31 36002 2.1 arc 21:33 36003 2.1 arc ----- sequence 36004, plate -9999 ------- 23:39 36004 0.0 bias 23:53 36005 0.0 bias 23:59 36006 0.0 bias ----- sequence 36007, plate 1542 ------- 25.8 20.1 29.2 17.8 DONE 01:05 36007 10.0 flat MJD 53733 17.4 12.0 19.9 12.1 01:08 36008 2.0 arc lowz 01:37 36009 900.1 target 01:55 36010 900.1 target 02:12 36011 900.1 target 02:30 36012 900.1 target 02:33 36013 10.0 flat 02:35 36014 2.0 arc ----- sequence 36015, plate 1543 ------- 7.1 5.2 9.4 6.8 Continue. 02:49 36015 10.0 flat lowz 02:51 36016 2.0 arc 03:11 36017 900.1 target 03:28 36018 900.1 target ----- sequence 36019, plate 2067 ------- 26.4 22.9 25.8 19.2 Continue 03:40 36019 10.0 flat Seguefaint 03:43 36020 2.0 arc 04:02 36021 900.1 target 04:19 36022 900.1 target 04:37 36023 900.1 target 04:55 36024 900.1 target 05:12 36025 900.1 target ----- sequence 36026, plate 2295 ------- 16.6 18.2 17.5 17.3 DONE 07:36 36026 10.0 flat Chunk 07:38 36027 2.0 arc 07:57 36028 900.1 target 08:14 36029 900.1 target 08:35 36030 1080.1 target ----- sequence 36031, plate 2349 ------- 6.2 2.8 5.6 2.7 [MJD 53731] 08:51 36031 10.0 flat 6.5 3.8 6.2 4.1 [MJD 53733] 08:53 36032 2.0 arc 12.0 8.3 10.9 9.0 DONE 09:13 36033 900.1 target Chunk 09:33 36034 1080.1 target ----- sequence 36035, plate 2346 ------- 18.5 13.7 19.5 17.8 DONE 09:50 36035 10.0 flat 09:52 36036 2.0 arc 10:15 36037 1080.1 target 10:36 36038 1080.1 target 10:57 36039 1140.1 target ----- sequence 36040, plate 1975 ------- 30.0 18.6 31.8 16.7 DONE 11:14 36040 10.0 flat 11:16 36041 2.0 arc 11:42 36042 1200.1 target 12:15 36043 1800.1 target 12:47 36044 1800.1 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 01:20Z 2 1542 -13.78 -0.0019 56.45 0.0029 168.36 0.0129 1.000000 02:53Z 1 1543 -46.04 -0.0001 47.26 0.0022 142.64 -0.0312 1.000070 03:45Z 8 2067 -17.19 0.0001 47.30 0.0011 165.35 -0.0218 1.000070 07:40Z 5 2295 83.09 -0.0005 53.99 0.0041 244.30 -0.0121 1.000220 08:56Z 4 2349 79.95 0.0009 62.84 0.0035 245.41 0.0049 1.000180 09:55Z 9 2346 54.61 -0.0030 76.06 0.0024 228.59 0.0345 1.000180 11:22Z 3 1975 95.05 0.0008 66.96 0.0040 260.18 0.0165 1.000220 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 23:41Z 2 0 1.00000 0 1257 0 121 30.0 6.8 5 292 23:57Z 2 0 1.00000 0 1257 0 121 30.0 5.6 7 301 01:20Z 2 1542 1.00000 0 1282 120 -14 56.4 3.9 11 270 -- 1.3 02:54Z 1 1543 1.00007 -691 767 150 -46 47.2 2.9 14 254 -- 1.3 03:45Z 8 2067 1.00007 -691 767 150 -17 47.3 1.8 14 240 -- 1.4 07:40Z 5 2295 1.00022 -2173 -529 100 83 54.0 1.8 14 283 -- 1.8 08:56Z 4 2349 1.00018 -1778 -200 125 80 62.9 1.4 12 332 -- 1.8 09:55Z 9 2346 1.00018 -1778 -224 125 54 76.1 0.9 9 12 -- 2.1 11:22Z 3 1975 1.00022 -2173 -536 125 95 67.0 1.4 17 314 -- 3.0 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 21:06Z 42 17 8 356 (NNW) 93 - Mostly Clear 21:37Z 42 12 8 357 (NNW) 91 - Clear 22:07Z 42 11 12 336 (NNW) -9999 - Clear 22:38Z 45 9 13 292 (WNW) 89 - Clear 23:10Z 42 6 0 224 (SW) 125 - Clear 23:41Z 44 11 4 292 (WNW) 157 - Clear 00:12Z 40 8 3 270 (W) 161 - Clear 00:41Z 37 19 3 256 (WSW) 290 - 01:12Z 37 14 7 264 (W) 97 - 01:43Z 38 17 12 271 (W) 324 - 02:13Z 37 20 10 258 (WSW) 440 - 02:43Z 36 20 15 242 (WSW) 469 - 03:14Z 36 20 14 240 (WSW) 477 - 03:44Z 36 17 11 249 (WSW) 419 - 04:15Z 35 11 7 259 (W) 452 - 04:45Z 33 12 6 266 (W) 482 - 05:15Z 36 10 8 266 (W) 601 - 05:46Z 35 12 6 290 (WNW) 566 - Clouds 06:17Z 35 12 5 268 (W) 661 - 06:47Z 35 12 8 264 (W) 990 - 07:17Z 35 15 8 272 (W) 879 - 07:49Z 34 16 16 302 (WNW) 706 - 08:21Z 34 16 10 1 (N) 594 - 08:52Z 34 11 8 351 (NNW) 328 - 09:22Z 33 12 18 3 (N) 324 - 09:53Z 34 10 13 13 (NNE) 333 - 10:25Z 34 11 14 335 (NNW) 308 - 10:58Z 34 9 18 319 (NW) 305 - 11:29Z 34 7 14 325 (NW) 275 - 12:02Z 34 7 13 326 (NW) 315 - 12:34Z 33 5 5 356 (NNW) 241 - 13:05Z 34 -3 12 322 (NW) 203 - 13:30Z 36 -13 6 303 (WNW) 345 - ---=== TELESCOPE STATUS ===--- 23:10Z Doors Open, Louvers Open, Fans On 00:15Z Enclosure Open 13:15Z Fans off, louvers and doors closed, enclosure on Telescope parked at 121,30 with Cart 1 (Plate 1543) mounted Counterweights at 248 Spectro autofill reconnected No interlocks bypassed Dewar weights at 13:22Z Imager: 169lbs 27psi Spectro: 158lbs 19psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_9_2a Watcher: v2_31_0 MCP: v5_26_0 TPM: tpm_v3_1_0 AstroDa: v15_10 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_06 SoS: v5_1_0 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.8360 -5.8440 Axial B -5.6790 -5.6770 Axial C 0.9400 0.9420 Trans D -10.0030 -9.9940 Lateral E 10.3505 10.3759 Lateral F 11.9507 11.9380 GALILS Commanded: 4700. -2450. 3000. -7800. 3300. 3000. Actual: 4711. -2457. 2997. -7825. 3296. 3024. SETMIR VALUES PriDesOrient: 0.00 -1.68 7.05 385.25 92.02 PriOrient: 0.00 -1.52 7.02 384.99 93.81 SECONDARY: ---------- Focus: 0.00 Air Temp.: 6.0 Alt.: 29.997523 MIGS TONIGHT NOMINAL Axial A 1.5400 1.5420 Axial B 1.0560 1.0470 Axial C 1.1720 1.1650 Trans D -0.1150 -0.1240 GALILS Commanded: 1593488. 1603800. 1554219. -6800. -5900. Actual: 1593447. 1603642. 1554316. -6820. -5882. SETMIR VALUES SecDesOrient: 1257.00 15.87 0.00 -50.00 136.76 SecOrient: 1257.03 15.94 0.02 -49.54 136.69 ---=== PROBLEMS IN DETAIL ===--- Spectro Focus Details: sp1 sp2 Coll_motor_A 6000 -4490 Coll_motor_B 6000 -7770 Coll_motor_C 6000 -7100 im> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg 5.400 tpm_TM_TelairAvg.VALA 5.878 b1 r1 b2 r2 Mean Offset .40 -.11 .03 -.04 Recommended Piston 1012 407 Indicated best focus temperature between 3 and 4 C Note that when collimate is run in version v5_1_0 there is a warning message "speclog_version: Command not found." It doesn't appear to effect the calculations but is not seen in v4_10_7.